OpenStudy (anonymous):

Anyone good with Quadratic Formulas? I have a couple of questions for you. x^2+9x-4=0 -4x^2+5x-3=0

hero (hero):

You don't have to use quad formula to figure these out. You can complete the square. But if you want to use quadratic formula, here it is: \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a} \]

OpenStudy (anonymous):

Bacon

OpenStudy (phi):

to use the formula you should match your equation to a x^2 +b x + c=0 can you find a, b and c for the first equation?

OpenStudy (anonymous):

Im sorry, i truly don't understand